How many nanoseconds is this 5 x 10-9 s? Express these non-exponentially, and chose the unit...

  1. How many nanoseconds is this 5 x 10-9 s?
  1. Express these non-exponentially, and chose the unit and prefix that works best: 2 x 10-3 s; 1.5 x 10-4 l; 4 x 10-8 g
  1. If the normal resting cardiac output is around 5 liters per minute, how many ml is that in a second? Express in real numbers and exponentially.
  1. The rate of DNA replication is about 2 kb (kilobases) per minute. How many bases is that per second?
  1. Dr. Russo is filling out a prescription of allergy medication for a patient. The patient is to take 120 mg of medication a day for ninety days. If the pills are 15mg each, how many pills will be prescribed?
  1. Professor Larson is conducting research on branching morphogenesis in developing organs and glands. If a sample of salivary gland tissue develops cells at a rate of 12 cells every second, how many cells will the sample have grown in an hour?
  1. A textbook required for biology is listed in the bookstore for $162.60. You find out that the bookstore makes 38% of profit on that book. How much was the book the publisher sold the bookstore? Round to the nearest cent.
  1. The ABio202Z syllabus states that you need 89.5% of all points for an A-. There is a total of 600 points to achieve and there are two 2 point bonus assignments at the end of the semester, which you decide to skip. In the end, you achieve 534 points. What is your grade? What would it be if you had not skipped the bonus assignments?
  1. The ABIO202Z syllabus states that there are a total of 600 points, 120 of them associated with a final exam. If you have an average of 90% of points everything but the final exam, and you get only 50 points on the final exam, can you still get a B+? What is your final grade?
  1. If a typical Homo sapiens has a total of 8,000 white blood cells/ml of blood and 70% of them are neutrophils, how many neutrophils does he/she have?
  1. As of 2018, 8.3 billion metric tons (a billion tons is 109 tons)of plastic waste have been produced worldwide. If 0.747 billion metric tons of that is being recycled, what percent of waste is not recycled?
  1. Professor Mayerhofer is teaching a lecture on human biology and nutrition. A healthy individual has approximately 22% of their body weight as fat, and 35% as muscle. If an individual weighs 175lb, how much of that weight is fat and how much muscle?
  1. You have made a NaOH solution for an experiment with amylase in Professor Thurman’s class and your class mate is asking what the Molarity of it was. Your detailed lab notes say that you weighed out 3 g of NaOH and dissolved them in a total of 150 ml, but you did not write down the molarity. What was it? (MW NaOH 39.997g/mol). Round to 2 decimal places.
  1. Two solutions of sodium hydroxide are in storage at a lab. One has a concentration of hydrogen ions of 3.5x10-11, while the other has a concentration of 7.8x10-10. What are their pHs and which solution is more basic?
  1. You made a 0.3% KCl solution in Professor Forni’s lab but need to determine the Molarity of it for a paper. (MW of KCl is 74.55 g/mol)
Answer #1

Prefixes:
10^-3 = Milli
10^-6 = Micro
10^-9 = Nano

1.
1 nano second = 10^-9 second
5 X 10^-9 sec = 5 nanoseconds

2.
2 X 10^-3 sec = 2 millisec
1.5 X 10^-4 L = 150 X 10^-6 L = 150 microliter
4 X 10^-8 g = 0.4 X 10^-9 g = 0.4 nanograms

3.
1 L = 1000 mL
1 min = 60 sec
5 L/min = 5 X (1000/60) mL/sec = 83.33 mL/sec = 0.833 X 10^-2 mL/sec

4.
1 kb = 1000 bp
1 min = 60 sec
2 kb/min = 2 X (1000/60) bp/sec = 33.33 bp/sec

5. Required medication size = 120 mg
1 pill = 15 mg
Number of days = 9
Number of pills required per day = 120/15 = 8 pills
= 8 X 9 = 72 pills for 9 days